The cost p, in dollars, of having a dinner party catered by Elain's Kosher Catering Company can be approximated by the functon p(d)=60=15p, where p is the number of people attending the party. The banquet hall where the dinner wil be located can hold no more than 150 people. Find the domain of the function.
Answer by checkley77(12844)

p(d)=60=15p
One too many = signs.
Assuming you mean:
p(d)=60+15p
Then:
p(150)=60+15*150
p(150)=60+2,250
p(150)=$2,310 for a party of 150.